OpenStudy (anonymous):

Consider an object moving in the plane whose location at time t seconds is given by the parametric equations: x(t)=4cos(πt) y(t)=3sin(πt). Assume the distance units in the plane are meters. (x^2/a^2)+(y^2/b^2)=1 What are a and b?

Parth (parthkohli):

So think of it like this. What would you do with these equations to eliminate the sine and cosine? Here's a hint:\[x = \cos(\pi t)\]\[y = \sin (\pi t)\]\[\Rightarrow x^2 + y^2 = \cos^2 (\pi t) + \sin^2 (\pi t) \tag{square both equations and add}\]But \(\cos^2 (\pi t) +\sin^2 (\pi t) = 1\), so\[x^2 + y^2 = 1\]

Parth (parthkohli):

It's just the same strategy except here, you've got \(x/4 = \cos(\pi t)\) and \(y/3 = \sin(\pi t)\).

OpenStudy (anonymous):

Right so after solving for y and x I just plug back in to the equation with a and b and solve for those?

OpenStudy (mathmale):

Can't solve for y and x. Rather, follow Parth Kohli's approach, which uses the Pythagorean Theorem to obtain the equivalent of x^2 + y^2 = 1 and thus eliminates the parameter t. Show your work, please. Hint: What geometric figure is represented by\[\frac{ x^2 }{ a^2}+\frac{ y^2 }{ b^2 }=1?\]

OpenStudy (anonymous):

An elipse is shown

Parth (parthkohli):

Solve for cosine and sine, actually, then square and add.

OpenStudy (anonymous):

I've gotten where I have 1=(x/4+y/3)^2